OpenStudy (anonymous):

HELP!!!!!!! WILL FAN AND MEDAL (im like really bad at this, so if you want to explain, explain it like if i was a 10 year old lol, but if you dont want to explain, you can just give me the answer) ------------------ Which statements are true, if f(x) = x + 1 and g(x)= (x – 1)^2? Check all that apply. 1) For any value of x, f(x) > g(x). 2) For any value of x, f(x) < g(x). 3) When x = 4, f(x) < g(x). 4) When x = 2, f(x) < g(x). 5) For all positive values of x, f(x) < g(x). 6) For all negative values of x, f(x) < g(x).

OpenStudy (tbnrfrags):

yeah im to stupid

OpenStudy (anonymous):

@ganeshie8 @cwrw238 @PaulaLovesSchool13

OpenStudy (anonymous):

Of course ^_^

OpenStudy (anonymous):

StatCounter - Free Web Tracker and Counter TheMathPage Topics in P R E C A L C U L U S Table of Contents | Home Share on facebookShare on twitterShare on printMore Sharing Services83 19 INVERSE FUNCTIONS Definition of inverses Extracting the argument Writing the inverse Notation The graph of an inverse function THE INVERSE of a function reverses the action of that function. Inverse function Say, for example, that a function f acts on 5, producing f(5). Then if g is the inverse of f, then g acting on f(5) will bring back 5. g(f(5)) = 5. Actually, g must do that for all values in the domain of f. And f must do that for all values in the domain of g. Here is the definition: Functions f(x) and g(x) are inverses of one another if: f(g(x)) = x and g(f(x)) = x, for all values of x in their respective domains. Problem 1. Let f(x) and g(x) be inverses. Then if f(0) = 8, what is the value of g(8)? To see the answer, pass your mouse over the colored area. To cover the answer again, click "Refresh" ("Reload"). g(8) = 0. For, f, acting on 0, produces 8. Therefore, since g is the inverse, then when it acts on 8, it will bring back 0. g(f(x)) = x. Example 1. Addition and subtraction are inverses. Subtracting a specific number reverses, or undoes, the result of adding it. In the language of functions, let f(x) = x + 2, and g(x) = x − 2. f(x) adds 2 to its argument. g(x) subtracts 2. Upon applying the definition: f(g(x)) = f(x − 2) = (x − 2) + 2 = x, and g(f(x)) = g(x + 2) = (x + 2) − 2 = x. Topic 3 The definition is satisfied. The functions f and g are inverses. Problem 2. Let f(x) = x2 and g(x) = x½. Show that they are inverses of one another. (The domain of f must be restricted to x greater than or equal to 0.) f(g(x)) = f(x½) = (x½)2 = x, and g(f(x)) = g(x2) = (x2)½ = x. Extracting the argument When we write (x + 3)4, then x + 3 is the argument of the function f(x) = x4. f is that function which takes the 4th power of its argument. Its inverse, g(x), will take the 4th root. g(x) = x¼. Example 2. Solve for x: (x + 3)4 = 16. Solution. To do that, we must free, or extract, the argument x + 3. We must write x + 3 = . . . And to extract the argument of any function, simply take its inverse. In this example, we take the 4th root of both sides of the equation. We can immediately write x + 3 = 16¼ = 2. Therefore, x = 2 − 3 = −1. Problem 3. Solve for x: Inverse function The inverse of taking the 5th root is taking the 5th power. Therefore, on taking the 5th power of both sides -- and thus freeing the argument: x − 4 = 25 = 32. x = 36. Writing the inverse The inverse of any function should be immediately clear. The inverse of x + 2 is x − 2. (Example 1) The inverse of x2 is x½. (Problem 2.) In Topic 21 we will see that the inverse of y = logbx is the exponential function y = bx. And in Topic 19 of Trigonometry, we see that the inverse of y = sin x is y = arcsin x. But say that we want to write the inverse of a this function: y = 3x − 4. Then we can invert it by solving for x. At each step of the solution, we apply an inverse. Upon exchanging sides: Inverse function We went from line (1) to line (2) because the inverse of subtracting 4 is adding 4. And we went from line (2) to line (3) because the inverse of multiplying by 3 is dividing by 3. If we now exchange the variables -- y = x 3 + 4 3 . -- then that function is the inverse of y = 3x − 4. In other words, the inverse of the function that first multiplies by 3 and then subtracts 4 -- y = 3x − 4 -- is the function that first adds 4 -- y = x + 4 -- and then divides by 3: y = x 3 + 4 3 . Problem 4. a) Write the inverse of f(x) = −5x. g(x) = x −5 = − x 5 Dividing by −5 is the inverse of multiplying by −5. b) Prove that they are inverses. f(g(x)) = f(− x 5 ) = −5· x −5 = x. And g(f(x)) = g(−5x) = −5x −5 = x. f multiplies its argument by −5. g divides its argument by −5. Problem 5. a) Let f(x) = −½x + 1. Can you immediately write g(x) its inverse? g(x) = −2x + 2. For, f is the function that multiplies its argument by −½ --equivalently, divides by −2 -- and then adds 1. Its inverse will therefore first subtract 1: x − 1 and then multiply by −2: −2(x − 1) = −2x + 2. b) Prove that f(x) and g(x) are inverses. f(g(x)) = −½(−2x + 2) + 1 = x − 1 + 1 = x. and g(f(x)) = −2(−½x + 1) + 2 = x − 2 + 2 = x. One sometimes sees that to "find" the inverse of a function, it is necessary to solve for x and exchange the variables. That never comes up in calculus. In fact, to solve for x in calculus requires knowing the inverse. (If y = x2, then to solve for x, one must know that the inverse is x½. Compare Problem 2.) And so the examples and problems given here are purely academic. They never come up in practice; they are simply examples of what inverse functions are. Notation The function I(x) = x is called the identity function. It always returns x. As a notation for the inverse of a function f, we sometimes see f −1 ("f inverse"). "−1" is not an exponent. That notation is used because in the language of composition of functions, we can write: f o f −1 = I This is similar in form to the multiplication of numbers, a· a−1 = 1. * * * For the inverse trigonometric functions, see Topic 19 of Trigonometry. The graph of an inverse function The graph of the inverse of a function f(x) can be found as follows: Inverse function Reflect the graph about the x-axis, then rotate it 90° counterclockwise (If we take the graph on the left to be the right-hand branch of y = x2, then the graph on the right is its inverse, y = Square root of x.) To see that that is the graph of the inverse, let A be any point on Inverse function the graph of f(x), let its coördinates be (a, b), let it be a distance d from the origin C, and let AC make an angle θ with the x-axis; triangle ABC is right angled. The figure on the left shows the reflection of A about the x-axis to the point D. The figure on the right shows the rotation of D 90° counterclockwise to the point C'. We will see that the coördinates of C' are (b, a) -- and those are coördinates on the graph of the inverse of f (x)Exclamation! For if we call that inverse g(x), then according to the figure on the left, f (a) = b. And g(b) -- the figure on the right -- returns us to a: g(b) = a. The definition of the inverse is satisfied. Inverse function To see that the coördinates of C' are (b, a), consider that since angle C'A'D is 90°, then C'A' makes an angle of 90° − θ with the x-axis. That is, angle C'A'B' is the complement of angle B'A'D, which is angle θ. Therefore in the right triangle A'B'C', the angle at C' is equal to θ. But the angle at A is the complement of θ. Therefore the triangles ABC, A'B'C' are congruent (Angle-side-angle), and those sides are equal that are opposite the equal angles: A'B' is equal to AB -- which is b, the y-coördinate of f (x). B'C' is equal to BC -- which is a, the x-coördinate of f (x). Therefore the coördinates of C' are (b, a). Inverse function So, when each point (a, b) on f(x) is transformed into (b, a), then the graph that results is its inverse. Each point (a, b) will also be transformed into (b, a) when (a, b) is reflected about the line y = x. Inverse function Therefore we say that the graphs of a function and its inverse are symmetrical with respect to the straight line y = x. End of the lessson Next Topic: Logarithms

OpenStudy (anonymous):

oh wait, i forgot the pic lol brb

OpenStudy (anonymous):

I think you should look here: http://www.themathpage.com/aprecalc/inverse-functions.htm

OpenStudy (anonymous):

Bubble look here: http://www.themathpage.com/aprecalc/inverse-functions.htm

OpenStudy (anonymous):

ok that looks like it was alot of help, thanks paula

OpenStudy (anonymous):

Sure Bubble :)

OpenStudy (anonymous):

Thanks for the medal ^_^

OpenStudy (anonymous):

np